OpenStudy (anonymous):

How can I show that the determinant of matrix A is the same as the determinant of the transpose of matrix A?

OpenStudy (anonymous):

\[\left( \begin{array}{cc} a & b \\ c & d \end{array} \right)\] ad-bc \[ \left( \begin{array}{cc} a & c \\ b & d \end{array} \right)\] ad-cb

OpenStudy (anonymous):

thanks but how can I show this fact for n-order of matrices?

OpenStudy (anonymous):

Induction...

OpenStudy (anonymous):

thanks.